OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 2x – 3y – z = –1 x – 4y + 3z = –11 2x + 9y + 3z = –13 a) –11y + 5z = 21 y + 9z = 9 b) 5y – 7z = 21 17y – 3z = 9 c) 7x – 13y = –14 8x + 0y = –16 d) –5x +5y = –8 –4x + 18z = –10

OpenStudy (amistre64):

what does the isolated equation look like when solved for x?

OpenStudy (amistre64):

c and d are out since they still retain an x ....so its gonna be a toss up between a and b i believe

OpenStudy (anonymous):

b is the answer multiply 2 to the second equation and subtract it from remaining equations

OpenStudy (cwrw238):

i think its a but the first equation dont seem correct

OpenStudy (amistre64):

it aint a, and i agree that its most likely b :)

OpenStudy (amistre64):

2*4 = 8 - 3 = 5y

OpenStudy (anonymous):

Thank you yall!

OpenStudy (cwrw238):

yea - its b

OpenStudy (cwrw238):

yw